OpenStudy (anonymous):

Marty's class collected $17.80 in a school-wide penny collection contest. If Tommy's class added $3.90 to the amount they collected, x, and then multiplied the sum by 2, they would have the same amount of money as Marty's class. 2(x + $3.90)= $17.80 How much money did Tommy's class collect? $5.00 $12.80 $3.90 $8.90

OpenStudy (anonymous):

x = 5.00 5.00+3.90 = 8.90 8.90 x 2 = 17.80

OpenStudy (anonymous):

i think 8.90 but not sure

OpenStudy (texaschic101):

2(x + 3.90) = 17.80 --- divide both sides by 2 x + 3.90 = 17.80/2 x + 3.90 = 8.90 -- subtract 3.90 from both sides x = 8.90 - 3.90 x = 5.00

OpenStudy (texaschic101):

or you could distribute the 2 thru the parenthesis and work it out that way....but either way, you should get the same answer